# Emergent symmetries in prethermal phases of periodically driven quantum systems

### v1

@@ IV Prethermal time crystals
The idea of a time crystalline phase of matter which breaks time-translational symmetry was first proposed in the context of equilibrium Hamiltonians in Ref. [174]. However, it was later shown that such a phase can not exist [175, 176, 177] in equilibrium. As noted in Ref. 177, such a phase can be diagnosed by an unequal-time correlation function of local operators $\hat{A}$ and $\hat{B}$: $C({\bf r-r^{\prime}},t)=\langle\hat{A}({\bf r},t)\hat{B}({\bf r^{\prime}},0)\rangle$, where the expectation is taken with respect to a many-body ground state. The presence of a time crystalline phase then constitutes a time-dependent long-time value of $C$ in the thermodynamic limit: $\lim_{V\to\infty,|{\bf r}-{\bf r^{\prime}}|\to\infty}C=f(t)$, where $f(t)$ is a periodic function of time. Ref. 177 explicitly showed that for local operators $\hat{A}$ and $\hat{B}$, $C$ can at best be a constant proving impossibility of such a phase in an equilibrium setting.
It was subsequently found that such phases can be realized in periodically driven quantum matter [90, 91, 92, 93, 94, 95, 96, 97]; they are dubbed as Floquet time crystals. A central problem encountered in realizing such phases constitutes drive-induced heating; Floquet system tend to undergo fast thermalization to a featureless infinite temperature steady state [7].
Several proposals for circumventing this problem was put forth. The first of them constitutes applying strong disorder leading to realization of a many-body localized phase which evades such thermalization [178, 179, 180, 181]; this allows one to stabilize a discrete time crystalline phase [90, 91, 92, 93, 94, 95, 97, 182]. The second is to use a domain-wall confinement mediated obstruction of rapid heating in an Ising chain [183]. The third, which we shall be mostly interested in, is to consider a large drive frequency which suppresses energy absorption leading to a large prethermal regime where a time crystalline phase may be stabilized [90, 91, 92, 93, 94, 96, 184, 185, 186, 187]. This phenomenon has been experimentally investigated using several platforms such as ion traps [188, 189, 190], quantum circuits [191], NMR systems [192, 193], superconducting qubits [194, 195] and quantum simulators[196]. It has also been theoretically studied in the context of open quantum systems [197, 198, 199].
The key difference of such a prethermal time crystal with other prethermal phenomena discussed so far in this review involves the central role played by micromotion in realization of time crystals. To understand this, we consider a generic unitary evolution with $H=H_{0}(t)+H_{1}(t)$ operator given, in the interaction picture,by
@@
$=$ $U_{0}(T,0)e^{-iH_{F}^{(1)}T/\hbar}=Xe^{-iH_{F}^{(1)}T/\hbar}$
where $X^{m}=I$ and $[X,H_{F}^{(1)}]=0$. Note that $m=1$ reproduces the standard results discussed in earlier sections; in this case, the stroboscopic dynamics is completely controlled by $H_{F}^{(1)}$ which, being local in the prethermal regime, can not exhibit time crystalline behavior [93, 177]. The commutation of $X$ with $H_{F}^{(1)}$ is most simply understood when $m=2$ ($X^{2}=I$); in this case, the requirement $\tilde{U}(T,0)=\tilde{U}(2T,T)$ leads to the relation [93]
@@
from which the commutation $[X,H_{F}^{(1)}]=0$ follows (for a more rigorous proof, see Ref. 96). We note here that for $m\neq 1$, although $H_{F}^{(1)}$ is local in the prethermal regime, $\tilde{U}$ is not a local operator and thus the arguments of Ref. 177 can be circumvented in this case.
The commutation of $H_{F}^{(1)}$ and $X$, in the rotated frame, indicates the presence of an additional approximate emergent internal symmetry of $H_{F}^{(1)}$ which leads to time crystalline behavior; such a symmetry was clearly absent in the microscopic Hamiltonian $H(t)$. Due to this internal symmetry, $H_{F}^{(1)}$ can be thought to have $m$ symmetry sectors; in a frame which moves with $X$, the operators there evolve under $H_{F}^{(1)}$ within one such symmetry sector. A rotation back to the lab frame therefore induces oscillations of operators with period $mT$ leading to a subharmonic response at $\omega=\omega_{D}/m$. These oscillations persist up to a prethermal time $t^{\ast}$ beyond which higher order terms in the Floquet Hamiltonian spoils the internal symmetry leading to melting of the time crystal; for high drive frequency regime, $t^{\ast}\sim\exp[\omega_{D}]$ and the subharmonic oscillations persists for an exponentially long timescale.
To understand this phenomenon a bit more, we now consider a simple example which corresponds to a driven Ising chain with [96]
$H_{1}(t)$ $=$ $-J\sum_{\langle ij\rangle}\sigma_{i}^{z}\sigma_{j}^{z}-h_{z}\sum_{j}\sigma_{j}^{z}$
$H_{0}(t)$ $=$ $h_{x}(t)\sum_{j}\sigma_{j}^{x}$ (33)
The simplest realization of time crystalline behavior can be obtained by choosing a periodic kicked protocol at times $t=kT$, where $k\in Z$, yielding [96]
@@
This corresponds to $U_{0}(2T,0)=X^{2}=I$, where $X=\prod_{j}\sigma_{j}^{x}$ is a manifestly non-local operator. The corresponding $H_{F}$ can be computed within Floquet perturbation theory and the first order term, which controls the dynamics in the prethermal phase, is given by $H_{F}^{(1)}=-J\sum_{\langle ij\rangle}\sigma_{i}^{z}\sigma_{j}^{z}$ which commutes with $X$. The stroboscopic evolution operator is therefore given by (for this example ${\mathcal{U}}=I$ [186])
$U(T,0)$ $=$ $\left(\prod_{j}\sigma_{j}^{x}\right)e^{iTJ\sum_{\langle ij\rangle}\sigma_{i}^{z}\sigma_{j}^{z}/\hbar}$ (35)
@@
$M(nT)$ $=$ $\frac{1}{L}\sum_{j}\langle\psi(nT)|\sigma_{j}^{z}|\psi(nT)\rangle=(-1)^{n}M(0)$
@@
Interestingly the eigenstates $|E_{\pm}\rangle$ of $H_{F}^{(1)}$ with energy $E_{z}$ are macroscopic cat states, namely they are superposition of many-body Fock states given by $|z\rangle$ in the $\sigma^{z}$ basis and its time-reversed partner $|{\bar{z}}\rangle$: $|E_{\pm}\rangle=(|z\rangle\pm|{\bar{z}}\rangle)/\sqrt{2}$. These states satisfy $X|E_{\pm}\rangle=\pm|E_{\pm}\rangle$; in terms of these states one can write
@@
The subharmonic response can be easily worked out from such a description. We note that in the present scenario, it is easy to see from these consideration that an initial state with $M(0)=0$ will fail to show any oscillations; one requires an initial macroscopic polarization. This feature distinguishes prethermal time crystals from their MBL counterparts which shows stable oscillations irrespective of initial states [91].
@@ V Discussion
In this brief review, we have discussed several prethermal phenomena in Floquet quantum systems. The key feature behind most of these constitutes an additional approximate emergent symmetry which allows for the realization such phases. In all cases, such emergent symmetry constitutes commutation of an operator ${\hat{O}}$ with a first-order Floquet Hamiltonian $H_{F}^{(1)}$ computed using some approximation scheme in either high drive frequency or high drive amplitude regime; no such commutation exists for higher order terms in the Floquet Hamiltonian. This feature essentially leads to prethermal nature of such phases; they are destroyed at long times when these higher order terms become important. An exact computation of the prethermal timescale $t^{\ast}$ beyond which drive-induced heating takes over is a complicated problem; however, in high drive frequency regime, $t^{\ast}$ can be estimated and it turns out to vary exponentially in the drive frequency. Such a large $t^{\ast}$ makes these prethermal phenomena experimentally relevant.
@@
The most notable exception to this principle, for the problems addressed in this review, constitutes prethermal quantum scars. Some of these Floquet scars can be mapped to their equilibrium counterparts and can be understood in terms of emergent symmetries [84, 85]; however, there is another class of such scars which have no obvious equilibrium analogue and it seems to be difficult to understand their existence from the point of view of an emergent symmetry [86, 88]. This difficulty partly stems from either the fact that they are not eigenstates of $H_{F}^{(1)}$ and necessarily requires contribution from higher order terms in $H_{F}$ [86] or from the fact that they exist in a non-perturbative regime where analytic computation of $H_{F}$ is difficult [88].
@@
Another important aspect which is much less studied is the stability of the prethermal Floquet physics under experimentally relevant perturbations. The physical phenomena such as quantum scars, Hilbert space fragmentation, time crystalline order are comparatively new and mostly confined to fewer model Hamiltonians and fine tuned regimes, hence it will be worth exploring the stability of such physics against disorder and other symmetry breaking or restoring perturbations. To have an analytically solvable model for HSF, time crystals; similar to the AKLT model of exact quantum scars will enhance the understanding these phenomena to a great extent.
@@
It will be interesting to see these prethermal Floquet phases in higher dimensional system as they can be easily constructed in cold atom platforms. For example, Ref. 200 found exponentially many QMBS in 2D PXP model. It will be interesting to explore whether the Floquet version of this model or its generalization can also stabilize such scars [201] in higher dimension and in different lattice geometries. Similar open question remain in the case of strong Hilbert space fragmentation. Recently [121] studied a 2D ring exchange spin model in the presence of subsystem symmetries and found signatures of strong Hilbert space fragmentation; similar signatures were also found in Refs. 123, 124 in other models. It will be interesting to investigate if there are prethermal version of such higher-dimensional HSF in periodically driven models.
@@
A relatively less explored related area of research constitutes understanding the effect of external bath on these phenomena. Some work has been done in this direction for time crystals [197, 198, 199] and quantum scars in context of non-driven systems [168, 169]. However, for the latter class of problems, the existence or properties of Floquet scars have not been investigated in details. A similar generalization can be studied for systems with prethermal Hilbert space fragmentation. Such open or non-Hermitian systems could be useful in preventing heating through bath or non-Hermitian term induced dissipation; they also allow for non-trivial steady states which can be controlled by the corresponding Lindbladians. These features may lead to possibility of engineering steady states with imprints of prethermal effects of closed systems. This issue has been studied for driven non-hermitian Ising chains [60, 61]; it certainly deserves a wider study in broader context.
@@
In conclusion, we have discussed several aspects of prethermal phases of Floquet quantum matter from the perspective of an approximate emergent symmetry which is crucial to their existence. Such systems exhibits dynamic freezing and localization and Hilbert space fragmentation and hosts Floquet scars and time crystalline phases. We have also discussed some open questions which can be subject of future work in these areas.
@@ VI Acknowledgments
The authors thank Sayan Choudhury, Arnab Das, Somsubhra Ghosh, Bhaskar Mukherjee, Sourav Nandy, Mainak Pal, Indranil Paul, Diptiman Sen, and Arnab Sen for several discussions on related topics. KS thanks DST, India for support through SERB project JCB/2021/000030.

### v2

@@ IV Prethermal time crystals
The idea of a time crystalline phase of matter which breaks time-translational symmetry was first proposed in the context of equilibrium Hamiltonians in Ref. wil1 . However, it was later shown that such a phase can not exist bruno1 ; noz1 ; masaki1 in equilibrium. As noted in Ref. masaki1, , such a phase can be diagnosed by an unequal-time correlation function of local operators $\hat{A}$ and $\hat{B}$: $C({\bf r-r^{\prime}},t)=\langle\hat{A}({\bf r},t)\hat{B}({\bf r^{\prime}},0)\rangle$, where the expectation is taken with respect to a many-body ground state. The presence of a time crystalline phase then constitutes a time-dependent long-time value of $C$ in the thermodynamic limit:
$\lim_{V\to\infty,|{\bf r}-{\bf r^{\prime}}|\to\infty}C=f(t),$ (30)
where $f(t)$ is a periodic function of time. Ref. masaki1, explicitly showed that for local operators $\hat{A}$ and $\hat{B}$, $C$ can at best be a constant proving impossibility of such a phase in an equilibrium setting.
It was subsequently found that such phases can be realized in periodically driven quantum matter tcrev1 ; tcrev2 ; tcrev3 ; tcrev4 ; tcrev5 ; tcpap1 ; tcpap2 ; tcpap3 ; they are dubbed as Floquet time crystals. A central problem encountered in realizing such phases constitutes drive-induced heating; Floquet systems tend to undergo fast thermalization to a featureless infinite temperature steady state rev7 .
Several proposals for circumventing this problem were put forth. The first of them constitutes applying strong disorder leading to realization of a many-body localized phase which evades such thermalization huserev1 ; ponte1 ; laza1 ; aba1 ; this allows one to stabilize a discrete time crystalline phase tcrev1 ; tcrev2 ; tcrev3 ; tcrev4 ; tcrev5 ; tcpap1 ; tcpap3 ; sondhipap1 . The second is to use a domain-wall confinement mediated obstruction of rapid heating in an Ising chain domain1 . The third, which we shall be mostly interested in, is to consider a large drive frequency which suppresses energy absorption leading to a large prethermal regime where a time crystalline phase may be stabilized tcrev1 ; tcrev2 ; tcrev3 ; tcrev4 ; tcrev5 ; tcpap2 ; luitz1 ; sheng1 ; machado1 ; gam1 . This phenomenon has been experimentally investigated using several platforms such as ion traps tcexp1 ; tcexp2 ; tcexp3 , quantum circuits tcexp4 , NMR systems tcexp5 ; tcexp6 , superconducting qubits tcexp7 ; tcexp8 and quantum simulatorstcexp9 . It has also been theoretically studied in the context of open quantum systems tcueda1 ; tcdem1 ; tcdsarma1 .
The key difference of such a prethermal time crystal with other prethermal phenomena discussed so far in this review involves the central role played by micromotion in realization of time crystals. To understand this, we consider a generic unitary evolution operator $U(T,0)$ corresponding to $H=H_{0}(t)+H_{1}(t)$. In the interaction picture, $U(T,0)$ an be written as
@@
$=$ $U_{0}(T,0)e^{-iH_{F}^{(1)}T/\hbar}=Xe^{-iH_{F}^{(1)}T/\hbar}.$
Here $X^{m}=I$, $[X,H_{F}^{(1)}]=0$, and $H_{F}^{(1)}$ is the perturbative first-order Floquet Hamiltonian. Note that $m=1$ reproduces the standard results discussed in earlier sections; in this case, the stroboscopic dynamics is completely controlled by $H_{F}^{(1)}$ which, being local in the prethermal regime, can not exhibit time crystalline behavior tcrev4 ; masaki1 . The commutation of $X$ with $H_{F}^{(1)}$ is most simply understood when $m=2$ ($X^{2}=I$); in this case, the requirement $\tilde{U}(T,0)=\tilde{U}(2T,T)$ leads to the relation tcrev4
@@
from which the commutation $[X,H_{F}^{(1)}]=0$ follows (for a more rigorous proof, see Ref. tcpap2, ). We note here that for $m\neq 1$, although $H_{F}^{(1)}$ is local in the prethermal regime, $\tilde{U}$ is not a local operator. Thus the arguments of Ref. masaki1, can be circumvented in this case.
The commutation of $H_{F}^{(1)}$ and $X$, in the rotated frame, indicates the presence of an additional approximate emergent internal symmetry of $H_{F}^{(1)}$ which leads to time crystalline behavior. Such a symmetry was clearly absent in the microscopic Hamiltonian $H(t)$. Due to this internal symmetry, $H_{F}^{(1)}$ can be thought to have $m$ symmetry sectors. In a frame which moves with $X$, the operators there evolve under $H_{F}^{(1)}$ within one such symmetry sector.
A rotation back to the lab frame therefore induces oscillations of operators with period $mT$ leading to a subharmonic response at $\omega=\omega_{D}/m$. These oscillations persist up to a prethermal time $t^{\ast}$ beyond which higher order terms in the Floquet Hamiltonian spoils the internal symmetry leading to melting of the time crystal. For the high drive frequency regime, $t^{\ast}\sim\exp[\omega_{D}]$ and the subharmonic oscillations persists for an exponentially long timescale.
To understand this phenomenon a bit more, we now consider a simple example which corresponds to a driven Ising chain with tcpap2
$H_{1}(t)$ $=$ $-J\sum_{\langle ij\rangle}\sigma_{i}^{z}\sigma_{j}^{z}-h_{z}\sum_{j}\sigma_{j}^{z},$
$H_{0}(t)$ $=$ $h_{x}(t)\sum_{j}\sigma_{j}^{x}.$ (34)
The simplest realization of time crystalline behavior can be obtained by choosing a periodic kicked protocol at times $t=kT$, where $k\in Z$, yielding tcpap2
@@
This corresponds to $U_{0}(2T,0)=X^{2}=I$, where $X=\prod_{j}\sigma_{j}^{x}$ is a manifestly non-local operator. The corresponding $H_{F}$ can be computed within Floquet perturbation theory. The first order term, which controls the dynamics in the prethermal phase, is given by $H_{F}^{(1)}=-J\sum_{\langle ij\rangle}\sigma_{i}^{z}\sigma_{j}^{z}$. This term commutes with $X$. The stroboscopic evolution operator is therefore given by (for this example ${\mathcal{U}}=I$ machado1 )
$U(T,0)$ $=$ $\left(\prod_{j}\sigma_{j}^{x}\right)e^{iTJ\sum_{\langle ij\rangle}\sigma_{i}^{z}\sigma_{j}^{z}/\hbar}.$ (36)
@@
$M(nT)$ $=$ $\frac{1}{L}\sum_{j}\langle\psi(nT)|\sigma_{j}^{z}|\psi(nT)\rangle$ (37)
$=$ $(-1)^{n}M(0),$
@@
Interestingly the eigenstates $|E_{\pm}\rangle$ of $H_{F}^{(1)}$ with energy $E_{z}$ are macroscopic cat states. These are superposition of many-body Fock states given by $|z\rangle$ and its time-reversed partner $|{\bar{z}}\rangle$ with $|E_{\pm}\rangle=(|z\rangle\pm|{\bar{z}}\rangle)/\sqrt{2}$. These states satisfy $X|E_{\pm}\rangle=\pm|E_{\pm}\rangle$; in terms of these states one can write
@@
The subharmonic response can be easily worked out from such a description. We note that in the present scenario, it is easy to see from these considerations that an initial state with $M(0)=0$ will fail to show any oscillations; one requires an initial macroscopic polarization. This feature distinguishes prethermal time crystals from their MBL counterparts which show stable oscillations irrespective of the chosen initial state tcrev2 .
@@ V Discussion
In this brief review, we have discussed several prethermal phenomena in Floquet quantum systems. The key feature behind most of these constitutes an additional approximate emergent symmetry which allows for the realization such phenomena. In all cases, such emergent symmetry constitutes commutation of an operator ${\hat{O}}$ with a first-order Floquet Hamiltonian $H_{F}^{(1)}$ computed using some approximation scheme in either high drive frequency or high drive amplitude regime. No such commutation exists for higher order terms in the Floquet Hamiltonian. This feature essentially leads to prethermal nature of such phases; they are destroyed at long times when these higher order terms become important. An exact computation of the prethermal timescale $t^{\ast}$ beyond which drive-induced heating takes over is a complicated problem. However, in the high drive frequency regime, $t^{\ast}$ can be estimated; it turns out to vary exponentially with the drive frequency. Such a large $t^{\ast}$ makes these prethermal phenomena experimentally relevant.
@@
The most notable exception to this principle, for the problems addressed in this review, constitutes prethermal quantum scars. Some of these Floquet scars can be mapped to their equilibrium counterparts and can be understood in terms of emergent symmetries bm1 ; mituza1 . However, there is another class of such scars which have no obvious equilibrium analogue. It seems to be difficult to understand their existence from the point of view of an emergent symmetry bm2 ; papic1 . This difficulty partly stems from either the fact that they are not eigenstates of $H_{F}^{(1)}$ and necessarily require contribution from higher order terms in $H_{F}$ bm2 or from the fact that they exist in a non-perturbative regime where analytic computation of $H_{F}$ is difficult papic1 .
@@
Another important aspect which is much less studied is the stability of the prethermal Floquet physics under experimentally relevant perturbations. Several phenomena such as quantum scars, Hilbert space fragmentation, and time crystalline order are comparatively new. Their analysis, so far, is mostly confined to few model Hamiltonians and fine tuned regimes. The stability of such phenomena against disorder and other symmetry breaking or restoring perturbations will certainly be worth exploring. Similarly, an analytically tractable model for Hilbert-space fragmentation, similar to the AKLT model of exact quantum scars, will enhance the understanding these phenomena to a great extent and will be of interest.
@@
It will be interesting to see these prethermal Floquet phases in higher dimensional system as they can be easily constructed in cold atom platforms. For example, Ref. scar2d1, found exponentially many QMBS in 2D PXP model. It will be interesting to explore whether the Floquet version of this model or its generalization can also stabilize such scars scar2d2 in higher dimension and in different lattice geometries. Similar open question remain in the case of strong Hilbert space fragmentation. Recently as1 studied a 2D ring exchange spin model in the presence of subsystem symmetries and found signatures of strong Hilbert space fragmentation. Similar signatures were also found in Refs. frank2, ; sid1, in other models. It will be interesting to investigate if there are prethermal version of such higher-dimensional Hilbert-space fragmentation in periodically driven models.
@@
A relatively less explored related area of research constitutes understanding the effect of external bath on these phenomena. Some work has been done in this direction for time crystals tcueda1 ; tcdem1 ; tcdsarma1 and quantum scars in context of non-driven systems scaropen1 ; scaropen2 . However, for the latter class of problems, the existence or properties of Floquet scars have not been investigated in details. A similar generalization can be made for systems with prethermal Hilbert space fragmentation. Such open or non-Hermitian systems could be useful in preventing heating through bath or non-Hermitian term induced dissipation; they also allow for non-trivial steady states which can be controlled by the corresponding Lindbladians. These features may lead to possibility of engineering steady states with imprints of prethermal effects of closed systems. This issue has been studied for driven non-hermitian Ising chains tb1 ; tb2 ; it certainly deserves a wider study in a broader context.
@@
In conclusion, we have discussed several aspects of prethermal phases of Floquet quantum matter from the perspective of an approximate emergent symmetry which is crucial to their existence. Such systems exhibits dynamic freezing and localization and Hilbert space fragmentation. They also host Floquet scars and time crystalline phases. We have also discussed some open questions that can be subjects of future work in this area.
@@ VI Acknowledgments
The authors thank Sayan Choudhury, Arnab Das, Somsubhra Ghosh, Bhaskar Mukherjee, Sourav Nandy, Roopayan Ghosh, Mainak Pal, Indranil Paul, Arnab Sen and Diptiman Sen for several discussions on related topics. KS thanks DST, India for support through SERB project JCB/2021/000030.